# Exact Mobility Edges in a Disorder-Free Dimerized Stark Lattice with Effective Unbounded Hopping

### v2

@@ III Numerical Results
We corroborate our theoretical findings with numerical diagonalization of the Hamiltonian under open boundary conditions (OBCs), examining the convergence to the thermodynamic limit. OBCs are chosen to avoid the large potential discontinuity at the boundaries inherent to Stark systems under periodic boundary conditions. While OBCs discretize the spectrum and introduce localized edge modes, the bulk density of states and the energy position of the ME are thermodynamic properties insensitive to boundary conditions [Pastur_1992]. Unless otherwise stated, we use the set of parameters $t_{1}=1$, $t_{2}=1.2$, $F=0.6$, $N=1000$. To quantify the localization properties, we utilize the IPR and the associated fractal dimension $D_{2}$ [Thouless_1974, Evers_2008], defined as
@@
where $\sigma\in\{A,B\}$ denotes the sublattice index. For an extended state, $\text{IPR}\sim L^{-1}$ and $D_{2}\to 1$, whereas for a localized state, $\text{IPR}\sim\text{const}$ and $D_{2}\to 0$. We rely on these scaling indices rather than the Lyapunov exponent, as the latter is ill-defined for systems exhibiting hyper-exponential WS localization [Longhi_2023].
@@
Figure 3: Characterization of localization. (a) Fractal dimension ($D_{2}$) versus energy ($L=20001$, $N=10000$, truncated at $|E|<6t_{1}$). Black dashed lines mark the analytical mobility edges at $|E|=t_{2}$. (b) Finite-size scaling of inverse participation ratio (IPR) versus system size $L$ for representative states: extended continuum ($E\approx 0.5t_{2}$), unbounded branch ($m=50$), and bounded branch (m=10m=-10). Dashed line indicates the scaling $\text{IPR}\sim L^{-1.00}$. (c, d) Real-space probability density $|\psi_{n}|^{2}$ on A and B sublattices (log scale, truncated at $10^{-7}$) shows the spatial distribution for the representative states ($L=2001,N=1000$). Other parameters match those in Fig. 1 ($t_{1}=1$, $t_{2}=1.2$, and $F=0.6$).
@@
The global phase diagram is presented in Fig. 1(d). To examine the sharpness of the transition, we plot $D_{2}$ as a function of energy for a large system size ($N=10000$) for the whole spectrum in Fig. 3(a). A distinct step-like transition is observed at the analytical ME $|E|=t_{2}$. States within the continuum ($|E|<t_{2}$, red) exhibit $D_{2}\approx 1$, indicating delocalization. In the localized regime ($|E|>t_{2}$), two distinct behaviors appear: the unbounded branch ($m>0$, blue) shows $D_{2}\approx 0$ as expected, while the bounded branch ($m<0$, green) exhibits intermediate fractal dimensions ($0.2\lesssim D_{2}\lesssim 0.8$). This elevated $D_{2}$ arises from the branch’s accumulation near the critical point, making it difficult to distinguish from critical or extended states using a fixed system size.
@@
To resolve this ambiguity and address the broader challenge of distinguishing true extended states from finite-size artifacts (“pseudo-MEs”), we rely on finite-size scaling. As demonstrated in the context of mosaic Stark lattices [Longhi_2023], states that appear extended at moderate sizes ($L\sim 10^{4}$) may eventually localize at extremely large scales ($L>10^{7}$). To determine the behavior in the thermodynamic limit of both the continuum and the anomalous bounded branch, we perform a finite-size scaling analysis of the IPR up to $N=5\times 10^{8}$ (total unit cells $L=2N+1\approx 10^{9}$, total sites $2L\approx 2\times 10^{9}$). We provide more details about the large-scale calculations in Appendix D and the robustness check in Appendix E.
@@
Fig. 3(b) displays the IPR scaling for representative states from each branch. The extended continuum state (red) follows a power law $\text{IPR}\propto L^{-1.00}$ across six orders of magnitude without deviation, providing evidence that the delocalization is not a transient effect but a persistent feature of the thermodynamic limit. In contrast, the bounded branch (green), despite residing close to the ME, exhibits a saturation of IPR ($\text{slope}\approx 0$), consistent with its localized nature. The lower saturation value of the IPR for the bounded branch compared to the unbounded branch reflects a larger localization length, which is consistent with its proximity to the delocalization transition.
@@
Finally, we visualize the spatial structure of these states in Fig. 3(c) and (d). The wavefunction distributions reveal the distinct mechanisms governing each branch. The unbounded branch (blue) is highly concentrated on the B-sublattice, directly pinned by the Stark potential. The bounded branch (green) resides primarily on the A-sublattice, allowing it to minimize the potential energy penalty and spread over a larger (but finite) range. The extended state (red) is dominated by the A-sublattice, where it forms a non-decaying standing wave, while its weight on the B-sublattice is significantly suppressed, consistent with the asymptotic behavior predicted by our analytical derivation.
@@ IV Experimental Proposal and Stability
We propose an experimental realization of our model using photonic frequency synthetic dimensions on a TFLN platform. This approach leverages the high tunability of electro-optic modulation to engineer complex hopping phases and site-dependent potentials, as recently demonstrated for the Creutz ladder topology [PhotonicCreutz_2025] and Bloch oscillations [BlochOscillation_2016].
@@ IV.1 Setup
The experimental setup, illustrated in Fig. 4(a), consists of two coupled ring resonators, A and B, which represent the two legs of the ladder. The lattice sites correspond to the frequency modes of the resonators, spaced by the FSR. As proposed in Ref. [PhotonicCreutz_2025], the complex hoppings required for the Creutz ladder topology can be realized through a combination of intra-resonator and inter-resonator modulations. Specifically, electro-optic modulators (EOMs) on each resonator driven at the FSR frequency generate the horizontal hopping terms ($A_{n}\leftrightarrow A_{n+1},B_{n}\leftrightarrow B_{n+1}$) with phases $\pm it_{2}/2$ controlled by the radio frequency (RF) drive phase. The vertical hopping ($t_{1}$) and cross hopping ($t_{2}/2$) are realized by connecting the resonators via a tunable Mach-Zehnder Interferometer (MZI) modulated by both direct current (DC) and RF signals [PhotonicCreutz_2025].
Figure 4: Experimental proposal and robustness analysis. (a) Schematic of the proposed photonic platform. Two ring resonators (A and B) are coupled via a tunable Mach-Zehnder Interferometer (MZI). RF electrodes on the rings generate intra-resonator hopping ($\pm it_{2}/2$), while the MZI generates inter-resonator couplings ($t_{1},t_{2}/2$). A DC bias on resonator B detunes its FSR to tune the Stark potential gradient $F$. (b) Robustness of the fractal dimension ($D_{2}$) against leakage potential strength $\epsilon$ on the A-sublattice with $\epsilon/F\in\{0.001,0.01,0.1\}$ (system size $L=201,N=100$). The energy is truncated by $|E|<10$. (c, d) Simulated steady-state output spectra with leakage potential $\epsilon=0.01F$ and different loss rates $\kappa/t_{1}\in\{0,0.01,0.05\}$ from resonator A (c) and resonator B (d) under injection at the center mode of resonator A ($E=0$). The spectra are normalized to the maximum intensity at port A. Other parameters match those in Fig. 1 ($t_{1}=1$, $t_{2}=1.2$, and $F=0.6$).
The crucial ingredient of our model—a linear Stark potential acting exclusively on the B-sublattice—is implemented by detuning the FSR of resonator B relative to the modulation frequency. In the frame rotating at the modulation frequency $\omega_{\text{mod}}$, the effective on-site potential for mode $n$ is given by $V_{n}=n(\Omega_{\text{FSR}}-\omega_{\text{mod}})$ [BlochOscillation_2016]. By lithographically designing resonator B to have a slightly different FSR from resonator A ($\Omega_{B}\neq\Omega_{A}$) and locking the modulation frequency to resonator A ($\omega_{\text{mod}}=\Omega_{A}$), we achieve a flat potential on the A-sublattice ($V_{n}^{A}\approx 0$) and a linear gradient on the B-sublattice ($V_{n}^{B}=Fn$), where $F=\Omega_{B}-\Omega_{A}$. The FSR can be fine-tuned via a DC bias voltage on the EOM, allowing precise control of the Stark potential parameter. We provide more details on the derivation of effective Hamiltonian in photonic systems in Appendix F.1.
@@ IV.2 Robustness Analysis
First, the unavoidable mismatch in the central resonant frequencies of the two resonators introduces a constant energy offset $\delta_{BA}$ between the A and B sublattices, described by $H_{\text{offset}}=\sum_{n}\delta_{BA}b_{n}^{\dagger}b_{n}$. This term shifts the zero-energy reference of the B-leg [i.e. $\mathcal{V}(k,E)\to\mathcal{V}(k,E)+\delta_{BA}$ in Eq. (4)], adding a phase factor $\exp(-i\delta k/F)$ to the B-sublattice wavefunction in momentum space. Crucially, this constant shift does not alter the localization properties or the existence of the ME, as the quantization condition for discrete energies and the no-pole condition remains invariant.
Second, imperfect locking of the modulation frequency to the FSR of resonator A introduces a residual “leakage” potential $V_{n}^{A}=\epsilon n$. While any nonzero $\epsilon$ theoretically localizes the extended states in the thermodynamic limit, we must assess their stability within the finite bandwidth of realistic photonic lattices (several hundreds of modes [FrequencyComb_2019]). We simulate the fractal dimension ($D_{2}$) of the eigenstates for a system size $L=201$ ($N=100$) under a leakage potential $H_{\text{leak}}=\sum_{n}\epsilon na^{\dagger}_{n}a_{n}$. As shown in Fig. 4(b), the ME remains distinguishable and the continuum states retain their extended character ($D_{2}\approx 1$) for leakage ratios $\epsilon/F\lesssim 10^{-2}$. This indicates that frequency locking stability on the order of 1% of the Stark gradient is sufficient for observation.
@@
where $\mathbf{s}_{\text{in}}$ is localized at the injection site. We simulate the output spectrum for an injection energy $\Omega_{0}=0$ (center of the extended band) in Fig. 4(c,d). The results show that for low loss ($\kappa\leq 0.01t_{1}$), the spectrum exhibits a broad, flat distribution on the A-sublattice (theoretically the A-sublattice distribution dominates), characteristic of an extended state. However, as the loss increases to $\kappa=0.05t_{1}$, the spectrum becomes asymmetric and decays away from the injection site, limiting the effective propagation range. This highlights the necessity of the strong coupling regime ($\kappa\ll t_{1}$).
@@ IV.3 Feasibility Estimation
We adopt realistic parameters reported in state-of-the-art TFLN literature to perform a numerical feasibility study. High-Q resonators ($Q\sim 1.5\times 10^{6}$) with strong electro-optic coupling have been demonstrated [FrequencyComb_2019], achieving coupling strengths comparable to the FSR ($t_{1}\sim 0.6\times\text{FSR}$). For a typical FSR of $2\pi\times 10$ GHz, this yields a hopping strength $t_{1}\approx 2\pi\times 5$ GHz. Our model parameter $F=0.6t_{1}$ corresponds to a detuning of $\sim 2\pi\times 3$ GHz (30% of FSR). To avoid weakening the coupling efficiency due to large detuning [BlochOscillation_2016], we suggest operating with a smaller gradient (e.g., $F\sim 2\pi\times 1$ GHz, 10% of FSR), which remains well within the tunable range. Under this detuning, the $1\%$ potential leakage threshold requires a frequency locking accuracy of $2\pi\times 10$ MHz.
The primary challenge is the loss rate. An intrinsic Q-factor of $1.5\times 10^{6}$ at 1550 nm corresponds to an intrinsic decay rate $\kappa_{0}\approx 2\pi\times 130$ MHz. By operating in the under-coupled regime to minimize broadening from external coupling, the total linewidth $\kappa$ can be maintained close to this intrinsic limit. With $t_{1}\approx 2\pi\times 5$ GHz, the ratio $\kappa/t_{1}\approx 2.6\%$, which is slightly above our conservative threshold of 1%. Some potential approaches to further decrease this ratio include utilizing ultra-low loss TFLN fabrication techniques ($Q\sim 10^{7}$ [HighQ_2017, HighQ_2024]) or by increasing the FSR and RF drive efficiency [LowV_2019, EnhanceCoupling_2025] to boost the effective hopping strength. Thus, the observation of ME signatures is feasible with current or near-future TFLN technology.
@@ V Conclusion
In summary, we have proposed and solved a disorder-free dimerized Stark lattice hosting an exact mobility edge (ME). By mapping the system to an unbounded Jacobi matrix, we demonstrate that the model evades the restrictions of the Simon-Spencer theorem, enabling the existence of true extended states in the presence of an unbounded potential. Analytically, we identify the transition at $|E|=t_{2}$, separating a continuum of extended states from two distinct localized branches: a standard Wannier-Stark ladder and an anomalous bounded branch accumulating at the ME. These predictions are corroborated by large-scale finite-size scaling ($L\sim 10^{9}$), which supports the persistence of the extended phase in the thermodynamic limit.
Furthermore, we propose an experimental realization using photonic frequency synthetic dimensions on thin-film lithium niobate (TFLN) resonators. We perform a stability analysis simulating experimental imperfections based on realistic parameters. Our results suggest that the ME is robust against free spectral range (FSR) mismatch and photon loss, provided the system operates in the strong coupling regime. Our work not only resolves the theoretical ambiguity surrounding MEs in Stark systems but also establishes a general mechanism for realizing and observing delocalization in the presence of unbounded potentials using modern photonic platforms.

### v3

@@ III Numerical Results
We corroborate our theoretical findings with numerical diagonalization of the Hamiltonian under open boundary conditions (OBCs), examining the convergence to the thermodynamic limit. OBCs are chosen to avoid the large potential discontinuity at the boundaries inherent to Stark systems under periodic boundary conditions. While OBCs discretize the spectrum and introduce localized edge modes, the bulk density of states and the energy position of the ME are thermodynamic properties insensitive to boundary conditions [51]. Unless otherwise stated, we use the set of parameters $t_{1}=1$, $t_{2}=1.2$, $F=0.6$, $N=1000$. To quantify the localization properties, we utilize the IPR and the associated fractal dimension $D_{2}$ [67, 21], defined as
@@
where $\sigma\in\{A,B\}$ denotes the sublattice index. For an extended state, $\text{IPR}\sim L^{-1}$ and $D_{2}\to 1$, whereas for a localized state, $\text{IPR}\sim\text{const}$ and $D_{2}\to 0$. We rely on these scaling indices rather than the Lyapunov exponent, as the latter is ill-defined for systems exhibiting hyperexponential WS localization [45].
@@
Figure 3: Characterization of localization. (a) Fractal dimension ($D_{2}$) vs energy ($L=20001$, N=10  [^0] N=10\ 000, truncated at $|E|<6t_{1}$). Black dashed lines mark the analytical mobility edges at $|E|=t_{2}$. (b) Finite-size scaling of inverse participation ratio (IPR) vs system size $L$ for representative states: extended continuum ($E\approx 0.5t_{2}$), unbounded branch ($m=50$), and bounded branch (m=10m=-10). Dashed line indicates the scaling $\text{IPR}\sim L^{-1.00}$. (c),(d) Real-space probability density $|\psi_{n}|^{2}$ on AA and $B$ sublattices (logarithmic scale, truncated at $10^{-7}$) shows the spatial distribution for the representative states ($L=2001,N=1000$). Other parameters match those in Fig. 1 ($t_{1}=1$, $t_{2}=1.2$, and $F=0.6$).
@@
The global phase diagram is presented in Fig. 1(d). To examine the sharpness of the transition, we plot $D_{2}$ as a function of energy for a large system size (N=10  [^0] N=10\ 000) for the whole spectrum in Fig. 3(a). A distinct steplike transition is observed at the analytical ME $|E|=t_{2}$. States within the continuum ($|E|<t_{2}$, red) exhibit $D_{2}\approx 1$, indicating delocalization. In the localized regime ($|E|>t_{2}$), two distinct behaviors appear: the unbounded branch ($m>0$, blue) shows $D_{2}\approx 0$ as expected, while the bounded branch ($m<0$, green) exhibits intermediate fractal dimensions ($0.2\lesssim D_{2}\lesssim 0.8$). This elevated $D_{2}$ arises from the branch’s accumulation near the critical point, making it difficult to distinguish from critical or extended states using a fixed system size.
@@
To resolve this ambiguity and address the broader challenge of distinguishing true extended states from finite-size artifacts (“pseudo-MEs”), we rely on finite-size scaling. As demonstrated in the context of mosaic Stark lattices [45], states that appear extended at moderate sizes ($L\sim 10^{4}$) may eventually localize at extremely large scales ($L>10^{7}$). To determine the behavior in the thermodynamic limit of both the continuum and the anomalous bounded branch, we perform a finite-size scaling analysis of the IPR up to $N=5\times 10^{8}$ (total unit cells $L=2N+1\approx 10^{9}$, total sites $2L\approx 2\times 10^{9}$). We provide more details about the large-scale calculations in Appendix D and the robustness check in Appendix E.
@@
Figure 3(b) displays the IPR scaling for representative states from each branch. The extended continuum state (orange) follows a power law $\text{IPR}\propto L^{-1.00}$ across six orders of magnitude without deviation, providing evidence that the delocalization is not a transient effect but a persistent feature of the thermodynamic limit. In contrast, the bounded branch (green), despite residing close to the ME, exhibits a saturation of IPR ($\text{slope}\approx 0$), consistent with its localized nature. The lower saturation value of the IPR for the bounded branch compared to the unbounded branch reflects a larger localization length, which is consistent with its proximity to the delocalization transition.
@@
Finally, we visualize the spatial structure of these states in Figs. 3(c) and 3(d). The wavefunction distributions reveal the distinct mechanisms governing each branch. The unbounded branch (blue) is highly concentrated on the $B$ sublattice, directly pinned by the Stark potential. The bounded branch (green) resides primarily on the $A$ sublattice, allowing it to minimize the potential energy penalty and spread over a larger (but finite) range. The extended state (orange) is dominated by the $A$ sublattice, where it forms a nondecaying standing wave, while its weight on the $B$ sublattice is significantly suppressed, consistent with the asymptotic behavior predicted by our analytical derivation.
@@ IV Experimental Proposal and Stability
We propose an experimental realization of our model using photonic frequency synthetic dimensions on a TFLN platform. This approach leverages the high tunability of electro-optic modulation to engineer complex hopping phases and site-dependent potentials, as recently demonstrated for the Creutz ladder topology [74] and Bloch oscillations [79].
@@ IV.1 Setup
The experimental setup, illustrated in Fig. 4(a), consists of two coupled ring resonators, A and B, which represent the two legs of the ladder. The lattice sites correspond to the frequency modes of the resonators, spaced by the FSR. As proposed in Ref. [74], the complex hoppings required for the Creutz ladder topology can be realized through a combination of intraresonator and interresonator modulations. Specifically, electro-optic modulators (EOMs) on each resonator driven at the FSR frequency generate the horizontal hopping terms ($A_{n}\leftrightarrow A_{n+1},B_{n}\leftrightarrow B_{n+1}$) with phases $\pm it_{2}/2$ controlled by the radio frequency (rf) drive phase. The vertical hopping ($t_{1}$) and cross hopping ($t_{2}/2$) are realized by connecting the resonators via a tunable Mach-Zehnder Interferometer (MZI) modulated by both direct current (DC) and RF signals [74].
Figure 4: Experimental proposal and robustness analysis. (a) Schematic of the proposed photonic platform. Two ring resonators (A and B) are coupled via a tunable Mach-Zehnder Interferometer (MZI). RF electrodes on the rings generate intraresonator hopping ($\pm it_{2}/2$), while the MZI generates interresonator couplings ($t_{1},t_{2}/2$). A DC bias on resonator B detunes its FSR to tune the Stark potential gradient $F$. (b) Robustness of the fractal dimension ($D_{2}$) against leakage potential strength $\epsilon$ on the $A$ sublattice with $\epsilon/F\in\{0.001,0.01,0.1\}$ (system size $L=201,N=100$). The energy is truncated by $|E|<10$. (c, d) Simulated steady-state output spectra with leakage potential $\epsilon=0.01F$ and different loss rates $\kappa/t_{1}\in\{0,0.01,0.05\}$ from resonator A (c) and resonator B (d) under injection at the center mode of resonator A ($E=0$). The spectra are normalized to the maximum intensity at port A. Other parameters match those in Fig. 1 ($t_{1}=1$, $t_{2}=1.2$, and $F=0.6$).
The crucial ingredient of our model—a linear Stark potential acting exclusively on the $B$ sublattice—is implemented by detuning the FSR of resonator B relative to the modulation frequency. In the frame rotating at the modulation frequency $\omega_{\text{mod}}$, the effective on-site potential for mode $n$ is given by $V_{n}=n(\Omega_{\text{FSR}}-\omega_{\text{mod}})$ [79]. By lithographically designing resonator B to have a slightly different FSR from resonator A ($\Omega_{\text{B}}\neq\Omega_{\text{A}}$) and locking the modulation frequency to resonator A ($\omega_{\text{mod}}=\Omega_{\text{A}}$), we achieve a flat potential on the $A$ sublattice ($V_{n}^{\text{A}}\approx 0$) and a linear gradient on the $B$ sublattice ($V_{n}^{\text{B}}=Fn$), where $F=\Omega_{\text{B}}-\Omega_{\text{A}}$. The FSR can be fine-tuned via a DC bias voltage on the EOM, allowing precise control of the Stark potential parameter. We provide more details on the derivation of effective Hamiltonian in photonic systems in Appendix F.1.
@@ IV.2 Robustness analysis
First, the unavoidable mismatch in the central resonant frequencies of the two resonators introduces a constant energy offset $\delta_{\text{BA}}$ between the $A$ and $B$ sublattices, described by $H_{\text{offset}}=\sum_{n}\delta_{\text{BA}}b_{n}^{\dagger}b_{n}$. This term shifts the zero-energy reference of the $B$ leg [i.e. $\mathcal{V}(k,E)\to\mathcal{V}(k,E)+\delta_{\text{BA}}$ in Eq. (4)], adding a phase factor $\exp(-i\delta k/F)$ to the $B$ sublattice wavefunction in momentum space. Crucially, this constant shift does not alter the localization properties or the existence of the ME, as the quantization condition for discrete energies and the no-pole condition remains invariant.
Second, imperfect locking of the modulation frequency to the FSR of resonator A introduces a residual “leakage” potential $V_{n}^{\text{A}}=\epsilon n$. While any nonzero $\epsilon$ theoretically localizes the extended states in the thermodynamic limit, we must assess their stability within the finite bandwidth of realistic photonic lattices (several hundreds of modes [83]). We simulate the fractal dimension ($D_{2}$) of the eigenstates for a system size $L=201$ ($N=100$) under a leakage potential $H_{\text{leak}}=\sum_{n}\epsilon na^{\dagger}_{n}a_{n}$. As shown in Fig. 4(b), the ME remains distinguishable and the continuum states retain their extended character ($D_{2}\approx 1$) for leakage ratios $\epsilon/F\lesssim 10^{-2}$. This indicates that frequency locking stability on the order of 1% of the Stark gradient is sufficient for observation.
@@
where $\mathbf{s}_{\text{in}}$ is localized at the injection site. We simulate the output spectrum for an injection energy $\Omega_{0}=0$ (center of the extended band) in Figs. 4(c) and 4(d). The results show that for low loss ($\kappa\leq 0.01t_{1}$), the spectrum exhibits a broad, flat distribution on the $A$ sublattice (theoretically the $A$-sublattice distribution dominates), characteristic of an extended state. However, as the loss increases to $\kappa=0.05t_{1}$, the spectrum becomes asymmetric and decays away from the injection site, limiting the effective propagation range. This highlights the necessity of the strong coupling regime ($\kappa\ll t_{1}$).
@@ IV.3 Feasibility estimation
We adopt realistic parameters reported in state-of-the-art TFLN literature to perform a numerical feasibility study. High-$Q$ resonators ($Q\sim 1.5\times 10^{6}$) with strong electro-optic coupling have been demonstrated [83], achieving coupling strengths comparable to the FSR ($t_{1}\sim 0.6\times\text{FSR}$). For a typical FSR of $2\pi\times 10$ GHz, this yields a hopping strength $t_{1}\approx 2\pi\times 5$ GHz. Our model parameter $F=0.6t_{1}$ corresponds to a detuning of $\sim 2\pi\times 3$ GHz (30% of FSR). To avoid weakening the coupling efficiency due to large detuning [79], we suggest operating with a smaller gradient (e.g., $F\sim 2\pi\times 1$ GHz, 10% of FSR), which remains well within the tunable range. Under this detuning, the $1\%$ potential leakage threshold requires a frequency locking accuracy of $2\pi\times 10$ MHz.
The primary challenge is the loss rate. An intrinsic $Q$-factor of $1.5\times 10^{6}$ at 1550 nm corresponds to an intrinsic decay rate $\kappa_{0}\approx 2\pi\times 130$ MHz. By operating in the undercoupled regime to minimize broadening from external coupling, the total linewidth $\kappa$ can be maintained close to this intrinsic limit. With $t_{1}\approx 2\pi\times 5$ GHz, the ratio $\kappa/t_{1}\approx 2.6\%$, which is slightly above our conservative threshold of 1%. Some potential approaches to further decrease this ratio include utilizing ultralow loss TFLN fabrication techniques ($Q\sim 10^{7}$ [84, 90]) or by increasing the FSR and RF drive efficiency [54, 68] to boost the effective hopping strength. Thus, the observation of ME signatures is feasible with current or near-future TFLN technology.
@@ V Conclusion
In summary, we have proposed and solved a disorder-free dimerized Stark lattice hosting an exact ME. By mapping the system to an unbounded Jacobi matrix, we demonstrate that the model evades the restrictions of the Simon-Spencer theorem, enabling the existence of true extended states in the presence of an unbounded potential. Analytically, we identify the transition at $|E|=t_{2}$, separating a continuum of extended states from two distinct localized branches: a standard Wannier–Stark ladder and an anomalous bounded branch accumulating at the ME. These predictions are corroborated by large-scale finite-size scaling ($L\sim 10^{9}$), which supports the persistence of the extended phase in the thermodynamic limit.
Furthermore, we propose an experimental realization using photonic frequency synthetic dimensions on TFLN resonators. We perform a stability analysis simulating experimental imperfections based on realistic parameters. Our results suggest that the ME is robust against FSR mismatch and photon loss, provided the system operates in the strong coupling regime. Our work not only resolves the theoretical ambiguity surrounding MEs in Stark systems but also establishes a general mechanism for realizing and observing delocalization in the presence of unbounded potentials using modern photonic platforms.